Topstone Industries is expected to pay a dividend of $2.10 per share in one year. This dividend, along with firm’s earnings, etc., is expected to grow at a rate of 5% forever. If the current market price for a share of Topstone is $38.62 what is the cost of equity? (Answer: 10.44%)

steps required
Answer #1

Solution:

As per the Gordon growth model the price of a stock can be calculated using the following formula:

P0 = D1 / ( Ke – g )

Where,

P0 = Current Market Price of the stock    ;    D1 = Dividend payment in one year   ;

Ke = Cost of equity    ;   g = growth rate

As per the information given in the question we have

P0 = $ 38.62    ;   D1 = $ 2.10    ;    g = 5 % = 0.05   ; Ke = To find

Applying the above values in the formula we have

38.62 = 2.10 / ( Ke – 0.05 )

38.62 * ( Ke – 0.05 ) = 2.10

( Ke – 0.05 )= 2.10 / 38.62

( Ke – 0.05 ) = 0.054376

Ke = 0.05 + 0.054376

Ke = 0.104376

Ke = 10.4376 %

Ke = 10.44 %

Thus the Cost of Equity = 10.44 %